OpenStudy (anonymous):

Hao's dog ate 50 cans of dog food in 31 days. How many cans should Hao buy to feed his dog for the next 6 days

OpenStudy (tkhunny):

You need a unit rate. Calculate "Cans / Day"

OpenStudy (anonymous):

ik that but i need the answer.

OpenStudy (tkhunny):

Well, work it out. Let's see your best efforts. What do you get for "cans / day" based on the given information.

OpenStudy (anonymous):

40cans/31 days

OpenStudy (anonymous):

so... lemme thinkkk...

OpenStudy (tkhunny):

That's "canS / dayS". Now, calculate "cans / DAY" How many cans in just one day. This is the crux of the problem.

OpenStudy (anonymous):

well as i divide its a repeating decimal.

OpenStudy (tkhunny):

Sweet. So, what is it?

OpenStudy (anonymous):

do i estimate? (round)

OpenStudy (tkhunny):

Doesn't much matter, so long as you have a valid rate. Carry enough decimal places to make your result meaningful. If we're talking about cans of dog food over a week or so. 2-3 decimal places should be more than enough.

OpenStudy (anonymous):

So if I get 1.290322508........ and I just round to 1.29 is that good enough?

OpenStudy (tkhunny):

Can's of dog food don't come in ½ or ¼, so the task is to find the right INTEGER. That's cans / day. Multiply by the number of days and you're almost done.

OpenStudy (anonymous):

I learned about integers a loong time ago but i cant remember.

OpenStudy (tkhunny):

No worries. We can take a non-integer first. Then we'll worry about it.

OpenStudy (anonymous):

Can you just give me the answer please because i have 4 more hw packets and i really need to finish them

OpenStudy (tkhunny):

Already provided the method. Just do it. That's cans / day. Multiply by the number of days and you're almost done

OpenStudy (anonymous):

When you started talking abou integers i got confused so..

OpenStudy (tkhunny):

Okay, produce the answer without that concern and we can have that other conversation immediately following.

OpenStudy (anonymous):

So far i just have 7.74 as my answer.

OpenStudy (anonymous):

so do i round or do i keep the decimal

OpenStudy (tkhunny):

50/31 = 1.6129032258064516129032258064516 1.6129 will do. 6*(1.6129) = 9.6774 How did you get 7.74?

OpenStudy (anonymous):

I multiplied 1.29 6 times

OpenStudy (tkhunny):

I see. You're using 40/31. That's an error. It should be 50/31. Sorry I didn't see that earlier. Good work writing it down so I could find the error, though!

OpenStudy (anonymous):

oh the question is wrong i meant 40/31

OpenStudy (tkhunny):

Okay, then: 40/31 = 1.2903225806451612903225806451613 1.29 should do. 6*1.29 = 7.74 -- Just as you have it. Now for the integer. You CANNOT buy 7.74 cans. You must buy 8. You cannot get away with buying 7. Your dog will go hungry!! That would be cruel. That's it.

OpenStudy (anonymous):

thank you for your time i hope you didnt have aynything important

OpenStudy (tkhunny):

Note: 6*1 = 6 -- No good. 6*1.3 = 7.8 -- Okay 6*1.29 = 7.74 - Okay 6*1.2903 = 7.7418 -- Okay Rounding is important. If you round the RATE to Integer, your dog goes hungry. Just one decimal place was enough to cover this one for only 6 days.

OpenStudy (tkhunny):

I did have important things to do - helping you. That's why I'm here.